# Modeling Innovation Ecosystem Dynamics through Interacting Reinforced Bernoulli Processes

### v1

@@ Appendix C The case of a reducible interaction matrix
When the interaction matrix $\Gamma$ is not irreducible, it is possible to decompose it in irreducible sub-matrices such that the union of the spectra of the sub-matrices coincides with the spectrum of the original matrix. In the following, we will describe an heuristic argument (also employed in Iacopini et al. (2020) and Aletti et al. (2023a)), useful in order to detect the rate at which each $S_{t,h}$ grows along time in the case of a general matrix $\Gamma$.
@@
(C.1) ${\mathbf{S}}_{t}\approx{\mathbf{s}}(t)=\sum_{k=1}^{r}t^{\gamma_{k}}\sum_{i=0}^{p_{k}-1}\ln^{i}(t)\mathbf{c}_{i}$
@@ Appendix D Choice of the threshold (robustness check)
Figure 6. Behavior of the mean index per category over the years.
In this section we discuss the choice of the threshold $\tau$ that has been used in the data analysis of Section 4 to define when a given patent $n$ can be considered a success for (or in) category $h$, that is, if and only if $I_{n,h}>\tau$. Since we want to identify as a “success” only patents that have an extraordinary impact on at least one category, the value of the threshold $\tau$ should be greater than the index value of the vast majority of patents in the data set. However, this requirement is not so stringent in the sense that the percentage of patents with $(I_{n,h}>\tau)$ is already around $1\%$ for $\tau=0.1$ and is below $0.1\%$ for any $\tau>0.5$ (see Figure 6 and Table 3 for further details). This means that the majority of patents in the data set have an index value very low (precisely, below $0.1$). This is in accordance with Squicciarini et al. (2013), who observed that only a very small subset of patents typically receives a large number of forward citations and the mean value of the forward index decreases along time.
@@
Table 3. For each category $h$, percentage of patents with a value of the index $I_{n,h}$ greater than $\tau=0.1,0.3,0.5,0.7,0.9$.
Figure 7. Different slopes estimated through linear regression on $S_{t,h}$, in $\log_{10}-\log_{10}$ scale, for each category $h$ (color line) and each value of the threshold $\tau$ (x-axis). The solid black line indicates the common slope $\widehat{\gamma}^{*}$ estimated imposing the same slope for all categories. The dashed horizontal black line indicates the common slope $\widehat{\gamma}^{*}=0.689$ estimated for $\tau=0.8$ in Section 4.
@@
Figure 8. Goodness-of-fit index $R^{2}$ of the linear regression on $S_{t,h}$, in $\log_{10}-\log_{10}$ scale, for each value of the threshold $\tau$ (x-axis). The red line indicates the $R^{2}$ obtained imposing a common slope and the blue line the $R^{2}$ obtained allowing the slopes to be different across categories. Both the lines are always very close to each other and always higher than $0.95$ for any value of $\tau$.

### v2

@@ Appendix C The case of a reducible interaction matrix
When the interaction matrix $\Gamma$ is not irreducible, it is possible to decompose it in irreducible sub-matrices such that the union of the spectra of the sub-matrices coincides with the spectrum of the original matrix. In the following, we will describe an heuristic argument (also employed in Iacopini et al. (2020) and Aletti, Crimaldi and Ghiglietti (2023)), useful in order to detect the rate at which each $S_{t,h}$ grows along time in the case of a general matrix $\Gamma$.
@@
${\mathbf{S}}_{t}\approx{\mathbf{s}}(t)=\sum_{k=1}^{r}t^{\gamma_{k}}\sum_{i=0}^{p_{k}-1}\ln^{i}(t)\mathbf{c}_{i}$ (18)
@@ Appendix D Choice of the threshold (robustness check)
Figure 6: Behavior of the mean index per category over the years.
In this section we discuss the choice of the threshold $\tau$ that has been used in the data analysis of Section 4 to define when a given patent $n$ can be considered a success for (or in) category $h$, that is, if and only if $I_{n,h}>\tau$. Since we want to identify as a “success” only patents that have an extraordinary impact on at least one category, the value of the threshold $\tau$ should be greater than the index value of the vast majority of patents in the data set. However, this requirement is not so stringent in the sense that the percentage of patents with $(I_{n,h}>\tau)$ is already around $1\%$ for $\tau=0.1$ and is below $0.1\%$ for any $\tau>0.5$ (see Figure 6 and Table 3 for further details). This means that the majority of patents in the data set have an index value very low (precisely, below $0.1$). This is in accordance with Squicciarini, Dernis and Criscuolo (2013), who observed that only a very small subset of patents typically receives a large number of forward citations and the mean value of the forward index decreases along time.
@@
Table 3: For each category $h$, percentage of patents with a value of the index $I_{n,h}$ greater than $\tau=0.1,0.3,0.5,0.7,0.9$.
Figure 7: Different slopes estimated through linear regression on $S_{t,h}$, in $\log_{10}-\log_{10}$ scale, for each category $h$ (color line) and each value of the threshold $\tau$ (x-axis). The solid black line indicates the common slope $\widehat{\gamma}^{*}$ estimated imposing the same slope for all categories. The dashed horizontal black line indicates the common slope $\widehat{\gamma}^{*}=0.689$ estimated for $\tau=0.8$ in Section 4.
@@
Figure 8: Goodness-of-fit index $R^{2}$ of the linear regression on $S_{t,h}$, in $\log_{10}-\log_{10}$ scale, for each value of the threshold $\tau$ (x-axis). The red line indicates the $R^{2}$ obtained imposing a common slope and the blue line the $R^{2}$ obtained allowing the slopes to be different across categories. Both the lines are always very close to each other and always higher than $0.95$ for any value of $\tau$.
Acknowledgments
The work is partially supported by the European Union - NextGenerationEU through the Italian Ministry of University and Research under the National Recovery and Resilience Plan (PNRR) - M4C2 - Investment 1.3, title [Public Sector Indicators for Sustainability and Wellbeing (PUWELL)] - Program [Growing Resilient, INclusive and Sustainable (GRINS)] - PE18 - CUP J33C22002910001.
This research was supported by the “Resilienza Economica e Digitale” (RED) project (CUP D67G23000060001) funded by the Italian Ministry of University and Research (MUR) as “Department of Excellence” (Dipartimenti di Eccellenza 2023-2027, Ministerial Decree no. 230/2022)
Author contributions statement
All the authors contributed equally to the present work.